Projects A and B are mutually exclusive. Project A costs $10,000 and is expected to generate cash inflows of $4,000 for 4 years. Project B costs $10,000 and is expected to generate a single cash flow in year 4 of $20,000. The cost of capital is 17.5%. What is the NPV of the project you should accept?

Group of answer choices

$865.73

$492.49

$3,911.49

$354.94
Answer #1

A:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=4000/1.175+4000/1.175^2+4000/1.175^3+4000/1.175^4

=10865.73

NPV=Present value of inflows-Present value of outflows

=10865.73-10,000

=865.73(Approx)

B:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=20,000/1.175^4

=10492.49

NPV=Present value of inflows-Present value of outflows

=10492.49-10,000

=$492.49(Approx)

Hence since projects are mutually exclusive;project A must be selected having higher NPV.

Hence the correct option is:

$865.73(Approx).